# Strong Lensing Effect and Quasinormal Modes of Oscillations of Black Holes in $\boldsymbol{f(R,T)}$ Gravity Theory

### v1

@@ Abstract
In this work, we analyze the strong lensing phenomenon and quasinormal modes (QNMs) in the case of black holes (BHs) surrounded by fluids within the framework of $f(R,T)$ gravity, adopting a minimally coupled model of the theory. Our analysis is conducted for three surrounding fields corresponding to three different values of the parameter $\omega$ of the equations of state, each representing a unique class of BH solutions. A universal method developed by V. Bozza is employed for strong lensing analysis and the WKB approximation method to compute the QNMs of oscillation of the BHs. The influence of the model parameter $\beta$ on the deflection angle and associated lensing coefficients is analyzed. Our findings on lensing reveal that smaller values of $\beta$ cause photon divergence at larger impact parameters, with the results converging to the Schwarzschild limit as $\beta\to-16.7551$ for (for $\omega=0)$, $\beta\to-18.8495$ (for $\omega=1/3)$ and $\beta\to-13.7088$ (for $\omega=-2/3)$. Extending the analysis to the supermassive BH SgrA*, we examine the outermost Einstein rings, estimate three lensing observables: angular position $\vartheta_{\infty}$, angular separation $s$ and relative magnification $r_{\text{mag}}$ for the BHs. For a specific value of $\beta$, BHs with different field configurations exhibit substantial variations in their observable properties. The variation of amplitude and damping of QNMs with respect to the model parameter $\beta$ is analyzed for the BHs. We found that the $\beta$ parameter has a direct correlation with the amplitude and an inverse relation with the damping of the QNMs. Further, we use the time domain analysis to verify the results and found a good match between the two methods.
@@ I Introduction
Black holes (BHs) have gained significant attention in recent times. The two major milestones of modern physics that have propelled the growing interest in the field of BH physics are the discovery of gravitational waves (GWs) by the LIGO and Virgo team in 2015 [1, 2, 3, 4, 5, 6], and the first ever image of the BH M87 at the center of M87 galaxy, by the Event Horizon Telescope (EHT) group in 2019 [7, 8, 9, 10, 11, 12]. Since then a lot of efforts have been directed towards a better understanding of the physics of BHs. Though a lot remains to be uncovered about BHs, some insights have been achieved and it is expected that advanced detectors of the future will shed more light regarding these mysterious objects. BHs came out as a mathematical solution of Einstein field equations in the General Relativity (GR) framework in 1916, obtained by Karl Schwarzschild for the vacuum case. Ever since a number of papers have derived BH solutions in different spacetimes and different gravity theories (see Ref. [13, 14, 15] for a review).
GR is the most fundamental and robust theory of gravity thus far. GR has been backed by two most fascinating and crucial observations of recent times in addition to earlier supporting observations [16]. One is the detection of GWs and the other is the snap of the direct image of the BH at the center of galaxy M87 [17] as mentioned already. However, there are some issues with GR that need to be addressed so that a complete theory of gravity can be obtained. GR faces the normalization issue in the high energy regime [18]. GR can not explain the dark components in the Universe, especially it fails in the IR regime as it cannot explain the accelerated expansion of the Universe as recently confirmed [19, 20] and can not predict anything regarding the existence of dark matter [21]. It has compatibility issues with quantum mechanical phenomena [22]. These issues have motivated researchers to look for modified as well as alternative theories of gravity (ATGs) [17, 23, 24, 25]. Physicists modify either the matter-energy sector or the geometry part of the Eistein field equations in their efforts to mitigate the issues faced by GR. The former way of modification leads to the theories of dark energy and latter one leads to the development of modified theories of gravity (MTGs). Some of the most pursued MTGs are $f(R)$ gravity [26, 27, 28], $f(R,T$) gravity [29, 30], Rastall gravity [33, 34], $f(R,L_{m}$) gravity [35] and so on. On the other hand ATGs are based on fundamentally different geometrical approaches than the approach of GR. One of the most studied ATGs is the $f(Q)$ gravity [31, 32] theory.
@@
In the strong field regime, Darwin’s pioneering work on the interpretation of Einstein’s equations for orbits around the sun to determine whether particles or light rays captured by the sun is served as the groundwork for subsequent developments in this area [45]. Over the years, investigations in the strong field limit have gained significant attention due to their potential to extract information about BHs [43, 49, 50, 51, 52, 53, 55, 54, 56, 57, 59, 58, 60]. A wide range of work has been performed on the strong deflection lensing extensively for Schwarzschild BHs [61, 62, 64, 63, 65]. Virbhadra and his co-workers have made a significant contribution to the study of gravitational lensing by introducing the lens equation in the investigations of the strong field regime [65, 64]. Later they analyzed the formation and properties of relativistic images for a Schwarzschild BH in this context [66, 67]. Further investigations expanding to other static and symmetric BH spacetimes have been performed by many researchers [68, 69, 50, 70, 71, 51, 72, 85]. In fact, the past decade has witnessed a significant progress in studies on gravitational lensing including investigations for naked singularities [74, 75, 76], wormholes [77, 78, 79, 80], and other exotic entities [81, 82]. A plethora of parallel efforts have been made on strong field deflection in GR as well as in context of MTGs, revealing remarkable phenomena such as relativistic images and shadow formation [84, 85, 83, 86, 87, 88, 96, 89, 90, 91, 93, 94, 95, 92, 97]. However, when one considers the MTGs, the deflection angle encapsulates additional information about deviations from the Schwarzschild or Kerr metrics. By analyzing the deflection in such scenarios, one can constrain the parameters of MTGs and assess their viability in describing astrophysical observations.
@@
Detection of GWs has opened up a new window to look at the Universe and develop our understanding on how gravity works in strong gravity regimes. When a BH is perturbed by some field, it oscillates and releases the excess energy to go back to its steady state. This release of energy is in the form of spacetime ripples with very minute amplitudes and decaying as it travels through spacetime. These are nothing but quasinormal modes (QNMs) of oscillations [98, 99]. QNMs have not been detected yet but it is believed that future detectors of GWs will be able to observe them. The concept of QNMs theoretically emerged for the first time via the work of Vishveshwara in 1970 [100]. Later, in 1975, Chandrasekhar and Detweiler computed the QNMs for the first time [101]. Schrutz and Will computed the QNMs for the first time using the WKB approximation method [102]. Since then many numerical as well as analytical methods have been employed to compute the QNMs. Konoplya developed the WKB method up to the 13th order along with the Padé averaging technique [103]. Other methods such as the asymptotic iteration method (AIM), Bernstein spectral method, Leaver’s technique, etc. have been widely used in the literature to compute the QNMs (see Ref. [98] for a review). In Ref. [104], the authors computed QNMs of BHs in scalar-vector-tensor (SVT) gravity with a surrounding quintessence field. They also studied the thermodynamic and optical properties of the obtained BH solution. Ref. [105] discusses BHs in MTGs in both static and rotating configurations and studies their optical properties. Recently, the authors in Ref. [106] computed a class of Kiselev BH solutions in the framework of $f(R,T)$ gravity. In another work [107], the authors obtained different solutions of BHs considering various models of the $f(R,T)$ gravity framework and studied their properties. QNMs of BH solutions in the framework of the Hu-Sawicki model of $f(R)$ gravity have been studied in Ref. [17].
@@
Motivated by the factors mentioned above, our study aims to explore the strong limit lensing characteristics and QNMs of BHs within the framework of the $f(R,T)$ gravity. The recent formulation of a BH spacetime in this theoretical framework [106, 107] has sparked our interest in exploring gravitational bending effects as well as QNMs of the BH. In the strong field lensing investigation, the technique proposed by V. Bozza [43] is adopted, whereas for QNMs calculations the WKB approximation is employed.
@@
The rest of the paper is structured as follows: In Section II, we outline the theoretical framework of the study. In Section III, we focus on the derivation of the gravitational bending angle in the strong field limit in the BH spacetime. Additionally, this section includes the computation and analysis of lensing observables by considering the supermassive BH SgrA* as a $f(R,T)$ gravity BH. Then we compute QNMs of the BH in Section IV. Section V presents the time profile analysis of the BH. Finally, we present our concluding remark in Section VI.
@@ II $\boldsymbol{f(R,T)}$ Gravity Black Holes
[…]
@@
[…]
@@
where $f_{1}(R)=R$ and $f_{2}(T)=T$ are considered. In this study, the trace of the energy-momentum tensor $T$ is connected to the spherically symmetric Kiselev BHs. The energy-momentum tensor governing Kiselev BHs is effectively associated with an anisotropic fluid, whose components satisfy the following relations [112, 106]:
@@
where $\rho$ represents the energy density of the fluid and $\omega$ is the equation of state parameter. Choosing the matter Lagrangian density $L_{m}=P=-\,(p_{r}+2\,p_{t})/3$, with the radial and tangential pressures $p_{r}=-\,\rho$ and $p_{t}=(3\,\omega+1)/2$ respectively, calculated from the general form of the energy-momentum tensor $T_{\mu\nu}=-\,p_{t}g_{\mu\nu}+(\rho+p_{t})u_{\mu}u_{\nu}+(p_{r}-p_{t})v_{\mu}v_{\nu}$ of an anisotropic fluid, where $u_{\mu}$ and $v_{\mu}$ represent the radial four-vectors and the four-velocity vectors respectively [113], equation (5) can be rewritten as
[…]
Using these results the solution is obtained for the metric function (6) of spacetime in the $f(R,T)$ gravity framework with the model (7) as follows [107]:
@@
where $M$ represents the total mass of the BH enclosed by the surrounding fluid and $c_{2}$ is a constant which will be taken as unity throughout the study [112, 107]. It is evident from equation (11) that this BH solution is independent of the parameter $\alpha$. Notably, in our analysis, we are concerned with three choices of the parameter $\omega$ corresponding to dust field ($\omega=0$), radiation field ($\omega=1/3$) and quintessence field ($\omega=-2/3$) around the BH which yield three different BH solutions. In the absence of fluid, the solution (11) simplifies to the Schwarzschild BH solution, but simply $\omega\to 0$, $\omega\to 1/3$ or $\omega\to-2/3$ do not reduce the solution to a BH surrounded by dust, radiation or a quintessence field in the framework of GR due to the presence of parameter $\beta$. When $\beta\to 0$, $\omega\to 0$, the solution corresponds to the metric of a Schwarzschild BH with an effective mass $M_{eff}=2M-c_{2}$ [106]. Similarly, for $\beta\to 0$, $\omega\to 1/3$, the solution behaves as the Reissner-Nordström BH with effective charge $q^{2}_{eff}=c_{2}$ [106, 107].
@@
Figure 1: The horizon structure of minimally coupled $f(R,T)$ gravity BHs with different values of $\omega$ as well as for different values of model parameter $\beta$.
It also needs to outline some other key aspects of the considered BHs’ solution to ensure the completeness of the subsequent analyses. As a part, we plot the variation of the function $A(r)$ with respect to radial distance $r$ as shown in Fig. 1 to analyze the horizon structures of the BHs under consideration. It is done by the root analysis of equation $A(r)=0$ for different values of model parameter $\beta$ as the positive roots of the equation $A(r)=0$ correspond to the locations of horizons of the BHs. It is observed that our BHs’ solution (11) exhibits a fixed horizon at $r=1$ irrespective of values $\beta\geq 0$ or $\beta<0$. This fixed horizon remains unaffected by $\omega$ as well. Of course, it is not seen as the event horizon for the whole domain of parameter $\beta$ in all the classes of BHs. In the first representation with $\omega=0$, for $\beta\geq 0$, it is the event and the only horizon of the considered BH. In contrast to this, one may find mostly two distinct horizons for $\beta<0$ along with the horizon at $r=1$. For example, for small negative values of the parameter $\beta$ the first (Cauchy) horizon appears at $r<1$, whereas for large negative values of $\beta$ the second (event) horizon shifts to $r>1$. However, at sufficiently large negative values of $\beta$, the BH again admits only a single horizon at $r=1$. In the second plot of the figure where $\omega=1/3$, it is observed that the BH admits two horizons with one fixed event horizon at $r=1$ and the Cauchy horizon at $r<1$ for $\beta>0$, whereas $\beta<0$ pushes the event horizon to $r>1$ region by making a transition of $r=1$ as the Cauchy horizon. In addition, there exists only a single horizon at the common point $r=1$ for $\beta=0$ in this situation also. Moreover, similar to the first consideration, the BH has a single horizon for significantly large negative values of $\beta$. On the other hand, BH with the quintessence field as a surrounding possesses an event horizon in the region $r>1$ for smaller negative values of $\beta$ with Cauchy horizon at $r=1$. Otherwise it shows a single horizon at the fixed position as mentioned. Interestingly, with the Cauchy horizon at $r=1$, the event horizons of our BHs coincide with the Schwarzschild BH horizon at $r=2$ for $\beta=-16.7551$, $\beta=-18.8495$ and $\beta=-13.7088$ in the first ($\omega=0$), second ($\omega=1/3$) and third ($\omega=-2/3$) situations respectively. Moreover, it is examined numerically that the modified BHs under consideration exhibit non-asymptotically flat behavior for $\beta<-16.7551$ in the $\omega=0$ case, $\beta<-18.8495$ when $\omega=1/3$ and $\beta<-13.7088$ in the $\omega=-2/3$ scenario. Recognizing the importance of these critical values of $\beta$ on the geometry of the spacetime of the $f(R,T)$ BHs, we limit our analysis to $\beta\geq-16.7551$, $\beta\geq-18.8495$ and $\beta\geq-13.7088$ for the respective cases. It can be inferred that despite the event horizons at the same location, the spacetime geometry of the considered BHs differs fundamentally from that of the Schwarzschild BH due to the presence of the Cauchy horizon. Thus observing such a horizon structure scenario, we can remark that the structure of a minimally coupled $f(R,T)$ gravity BHs changes typically and reflects a significant change in the spacetime’s causal structure as the parameter $\beta$ varies.

### v2

@@ Abstract
In this work, we analyze the strong lensing phenomenon and quasinormal modes (QNMs) in the case of black holes (BHs) surrounded by fluids within the framework of $f(R,T)$ gravity, adopting a minimally coupled model of the theory. Our analysis is conducted for three surrounding fields corresponding to three different values of the parameter $\omega$ of the equations of state, each representing a unique class of BH solutions. A universal method developed by V. Bozza is employed for strong lensing analysis and the WKB approximation method to compute the QNMs of oscillation of the BHs. The influences of the model parameters $\beta$ and $c_{2}$ on the deflection angle and associated lensing coefficients are analyzed. Our findings on lensing reveal that smaller values of $\beta$ and $c_{2}$ cause photon divergence at larger impact parameters as well as the lensing results converge to the Schwarzschild limit. Extending the analysis to the supermassive BH Sgr A*, we examine the outermost Einstein rings, estimate three lensing observables: angular position $\vartheta_{\infty}$, angular separation $s$ and relative magnification $r_{\text{mag}}$ for the BHs. For a specific values of $\beta$ and $c_{2}$, BHs with different field configurations exhibit substantial variations in their observable properties. The variation of amplitude and damping of QNMs with respect to the model parameter $\beta$ and $c_{2}$ is analyzed for the BHs. We found that the $\beta$ parameter has a direct correlation with the amplitude and an inverse relation with the damping of the QNMs, while $c_{2}$ has direct correlation with amplitude as well as damping. Further, we use the time domain analysis to verify the results and found a good match between the two methods.
@@ I Introduction
General Relativity (GR) is the most fundamental and robust theory of gravity thus far. GR has been backed by two most fascinating and crucial observations of recent times in addition to earlier supporting observations [1]. One is the detection of gravitational waves (GWs) by the LIGO and Virgo team in 2015 [2, 3, 4, 5, 6, 7] and the other is the snap of the direct image of the black hole (BH) at the center of galaxy M87 [8] by the Event Horizon Telescope (EHT) group in 2019 [9, 10, 11, 12, 13, 14]. However, there are some issues with GR that need to be addressed so that a complete theory of gravity can be obtained. GR faces the normalization issue in the high energy regime [15]. GR can not explain the dark components in the Universe, especially it fails in the IR regime as it cannot explain the accelerated expansion of the Universe as recently confirmed [16, 17] and can not predict anything regarding the existence of dark matter [18]. It has compatibility issues with quantum mechanical phenomena [19]. These issues have motivated researchers to look for modified as well as alternative theories of gravity (ATGs) [8, 20, 21, 22]. Physicists modify either the matter-energy sector or the geometry part of the Einstein field equations in their efforts to mitigate the issues faced by GR. The former way of modification leads to the theories of dark energy and the latter one leads to the development of modified theories of gravity (MTGs). Some of the most pursued MTGs are $f(R)$ gravity [23, 24, 25], $f(R,T$) gravity [26, 27, 28, 29], Rastall gravity [30, 31], $f(R,L_{m})$ gravity [32] and so on. On the other hand ATGs are based on fundamentally different geometrical approaches than the approach of GR. One of the most studied ATGs is the $f(Q)$ gravity [33, 34] theory.
BHs have gained significant attention in recent times. The two major milestones of modern physics that have propelled the growing interest in the field of BH physics are the already mentioned detection of GWs and the photograph of the first ever image of the black hole M87 at the center of the M87 galaxy. Since then a lot of efforts have been directed towards a better understanding of this aspect of BH physics [35, 36, 37]. Though a lot remains to be uncovered about BHs, some insights have been achieved and it is expected that advanced detectors of the future will shed more light regarding these mysterious objects. BHs came out as a mathematical solution of Einstein field equations in the GR framework in 1916, obtained by Karl Schwarzschild for the vacuum case. Ever since a number of papers have derived BH solutions in different spacetimes and different gravity theories (see Ref. [38, 39, 40] for a review).
@@
In the strong field regime, Darwin’s pioneering work on the interpretation of Einstein’s equations for orbits around the sun to determine whether particles or light rays captured by the sun is served as the groundwork for subsequent developments in this area [50]. Over the years, investigations in the strong field limit have gained significant attention due to their potential to extract information about BHs [48, 54, 55, 56, 57, 58, 60, 59, 61, 62, 64, 63, 65]. A wide range of work has been performed on the strong deflection lensing extensively for Schwarzschild BHs [66, 67, 69, 68, 70]. Virbhadra and his co-workers have made a significant contribution to the study of gravitational lensing by introducing the lens equation in the investigations of the strong field regime [70, 69]. Later they analyzed the formation and properties of relativistic images for a Schwarzschild BH in this context [71, 72]. Further investigations expanding to other static and symmetric BH spacetimes have been performed by many researchers [73, 74, 55, 75, 76, 56, 77, 78]. In fact, the past decade has witnessed a significant progress in studies on gravitational lensing including investigations for naked singularities [79, 80, 81], wormholes [82, 83, 84, 85], and other exotic entities [86, 87]. A plethora of parallel efforts have been made on strong field deflection in GR as well as in context of MTGs, revealing remarkable phenomena such as relativistic images and shadow formation [96, 89, 78, 97, 88, 90, 91, 92, 102, 93, 94, 95, 99, 100, 101, 98, 103]. However, when one considers the MTGs, the deflection angle encapsulates additional information about deviations from the Schwarzschild or Kerr metrics. By analyzing the deflection in such scenarios, one can constrain the parameters of MTGs and assess their viability in describing astrophysical observations.
@@
Detection of GWs has opened up a new window to look at the Universe and develop our understanding on how gravity works in strong gravity regimes. When a BH is perturbed by some field, it oscillates and releases the excess energy to go back to its steady state. This release of energy is in the form of spacetime ripples with very minute amplitudes and decaying as it travels through spacetime. These are nothing but quasinormal modes (QNMs) of oscillations [104, 105]. The concept of QNMs theoretically emerged for the first time via the work of Vishveshwara in 1970 [108]. Later, in 1975, Chandrasekhar and Detweiler computed the QNMs for the first time [109]. Schrutz and Will computed the QNMs for the first time using the WKB approximation method [110]. Since then many numerical as well as analytical methods have been employed to compute the QNMs. Konoplya developed the WKB method up to the 13th order along with the Padé averaging technique [111]. Other methods such as the asymptotic iteration method (AIM), Bernstein spectral method, Leaver’s technique, etc. have been widely used in the literature to compute the QNMs (see Ref. [104] for a review). In Ref. [112], the authors computed QNMs of BHs in scalar-vector-tensor (SVT) gravity with a surrounding quintessence field. They also studied the thermodynamic and optical properties of the obtained BH solution. Ref. [113] discusses BHs in MTGs in both static and rotating configurations and studies their optical properties. Recently, the authors in [114] computed a class of Kiselev BH solutions in the framework of $f(R,T)$ gravity. In another work [115], the authors obtained different solutions of BHs considering various models of $f(R,T)$ gravity framework and studied its properties. QNMs of black hole solutions in the framework of Hu-Sawicki model of $f(R)$ gravity has been studied in Ref. [8].
An important observational as well as theoretical aspect with regard to QNMs is that they can be used to test the cosmic censorship conjecture, which basically states that a singularity must exist within an event horizon. The abnormal behaviour of QNMs like growing modes and the absence of a well-defined QNM spectrum can point towards the presence of naked singularities [116]. Another such aspect that is relevant to the QNMs is the testing of the No-Hair theorem which states that the BH has only three properties: mass, charge and spin (angular momentum). Thus, QNMs should depend only on these three conserved quantities. Any deviation of QNMs from this dependency can be an indication of the presence of ‘hairs’ [117]. It is observed that the QNMs for some values of model parameters are found to be indeterminate, meaning that QNMs do not exist. It may signal presence of singularities in the BH spacetime. However, to the best of our knowledge, QNMs have not been detected yet but it is believed that future detectors of GWs will be able to observe them. In this respect, it should be mentioned that Ref. [106] claims to have detected QNMs for $l=m=2$ and $l=m=3$ modes. It should be noted that such claims are subject to criticism and signal-to-noise (SNR) issues. Having said that, some works [106, 107] have claimed to detect some modes of QNMs but a more robust analysis is anticipated in the near future with advanced future detectors for more clarity. Readers can refer to Ref. [107] for an overview of this topic.
@@
Motivated by the factors mentioned above, our study aims to explore the strong limit lensing characteristics and QNMs of BHs within the framework of the $f(R,T)$ gravity. The recent formulation of a BH spacetime in this theoretical framework [114, 115] has sparked our interest in exploring gravitational bending effects as well as QNMs of the BH. In the strong field lensing investigation, the technique proposed by V. Bozza [48] is adopted, whereas for QNMs calculations 3rd order WKB approximation is employed.
@@
The rest of the paper is structured as follows: In Section II, we outline the theoretical framework of the study. In Section III, we focus on the derivation of the gravitational bending angle in the strong field limit in the BH spacetime. Additionally, this section includes the computation and analysis of lensing observables by considering the supermassive BH SgrA* as a $f(R,T)$ gravity BH. Constraining of the model parameters using the EHT observational data of the Einstein ring is presented in this section. Then we compute QNMs of the BH in Section IV. Section V presents the time profile analysis of the BH. Finally, we present our concluding remark in Section VI.
@@ II $\boldsymbol{f(R,T)}$ Gravity Black Holes
$T_{\mu\nu}=-\,\frac{2\,\delta(\sqrt{-\,g}\mathcal{L}_{m})}{\sqrt{-\,g}\ \delta g^{\mu\nu}}=-\,2\ \frac{\partial\mathcal{L}_{m}}{\partial g^{\mu\nu}}+g_{\mu\nu}\,\mathcal{L}_{m}.$ (2)
@@
$\theta_{\mu\nu}=g^{ab}\,\frac{\delta{T}_{ab}}{\delta{g^{\mu\nu}}}=-\,2\,T_{\mu\nu}+g_{\mu\nu}\,\mathcal{L}_{m}-2\,g^{ab}\frac{\partial^{2}{\mathcal{L}_{m}}}{\partial{g^{\mu\nu}}\partial{g^{ab}}}.$ (5)
Importantly, the inclusion of the trace $T$ of the energy-momentum tensor $T_{\mu\nu}$ in the gravitational action of $f(R,T)$ gravity theory results in some extra terms in the field equations of the theory. Because of this, it can be treated as a two fluid theory with the effective energy-momentum tensor $T_{\mu\nu}^{eff}=T_{\mu\nu}+T_{\mu\nu}^{I}$. Here $T_{\mu\nu}$ represents the energy-momentum tensor for ordinary matter field and $T_{\mu\nu}^{I}$ is the matter-curvature interaction term which may act as a new matter source and can be defined for a geometrically originated anisotropic matter distribution.
@@
where, $f_{1}(R)=R$  and $f_{2}(T)=T$. In this study, $T$ is chosen as the trace of the energy-momentum tensor connected to the spherically symmetric Kiselev BHs. The energy-momentum tensor governing Kiselev BHs is effectively associated with an anisotropic fluid, whose components satisfy the following correlations [122, 114]:
@@
where $\rho$ represents energy density of the fluid and $\omega$ is the parameter of equation of state. Choosing the matter Lagrangian density $L_{m}=P=(-1/3)(p_{r}+2\,p_{t})$,  with the radial and tangential pressures $p_{r}=-\rho$ and $p_{t}=\frac{1}{2}(3\,\omega+1)$ respectively, calculated from the general form of the energy-momentum tensor $T_{\mu\nu}=-p_{t}g_{\mu\nu}+(\rho+p_{t})u_{\mu}u_{\nu}+(p_{r}-p_{t})v_{\mu}v_{\nu}$ of an anisotropic fluid, where $u_{\mu}$ and $v_{\mu}$ represent the radial-four vectors and the four velocity vectors respectively [123], equation (5) can be rewritten as
$\theta_{\mu\nu}=-2T_{\mu\nu}-\frac{1}{3}(p_{r}+2p_{t})g_{\mu\nu}$ (10)
Using these results the solution obtained for metric functions of spacetime in $f(R,T)$ gravity framework for the model (7) is as follows [115]
@@
where $M$ represents the total mass of the BH enclosed by the surrounding fluid, $\omega$ is the equation of state parameter of the fluid and $c_{2}$ is a constant. It is evident from equation (11) that this BH solution is independent of the parameter $\alpha$. Notably, in our
@@
Refer to caption   Refer to caption   Refer to caption
@@
Figure 1: The horizon structure of minimally coupled $f(R,T)$ gravity BHs with different values of $\omega$ as well as for different values of model parameters $\beta$ and $c_{2}$, considering $M=1$ for all the cases.
analysis, we are concerned with three choices of the parameter $\omega$ corresponding to dust field ($\omega=0$), radiation field ($\omega=1/3$) and quintessence field ($\omega=-2/3$) around the BH which yield three different BH solutions. In the absence of fluid, the solution (11) simplifies to the Schwarzschild BH solution, but simply $\omega\to 0$, $\omega\to 1/3$ or $\omega\to-2/3$ do not reduce the solution to a BH surrounded by dust, radiation or a quintessence field in the framework of GR due to the presence of parameter $\beta$. When $\beta\to 0$, $\omega\to 0$, the solution corresponds to the metric of a Schwarzschild BH with an effective mass $M_{eff}=2M-c_{2}$ [114]. Similarly, for $\beta\to 0$, $\omega\to 1/3$, the solution behaves as the Reissner-Nordström BH with effective charge $q^{2}_{eff}=c_{2}$ [114, 115].
It is also crucial to outline some other key aspects of the considered BHs’ solution to ensure the completeness of the subsequent analyses. As a part, the metric function $A(r)$ is plotted with respect to radial distance $r$ as shown in Fig. 1 to analyze the horizon structure of the BHs under consideration. It is done by the root analysis of equation $A(r)=0$ for different values of model parameters $\beta$ and $c_{2}$ respectively, as the positive roots of the equation $A(r)=0$ correspond to the locations of horizons of the BHs. In Fig. 1, we have presented three plots in the upper row for $c_{2}=0.5$, and $\omega=0,1/3,-2/3$ respectively, each showcasing different values of $\beta$ along with the GR case. It is observed that the BHs can possess at most two horizons for all three cases while the number may be less depending on the values of the parameters. Similarly, in the lower row, we have plotted three corresponding graphs by changing $c_{2}$ for $\beta=-11.8$ and three aforementioned values of $\omega$ respectively along with the GR case as well. Significantly, here also it is seen that the maximum number of horizons that the BHs can possess does not exceed two at a fixed value of $\beta$. It is observed that corresponding to each surrounding field, for both $\beta$ and $c_{2}$ varying scenarios, the outer horizon moves away from the GR case with increasing values of $\beta$ and $c_{2}$ respectively. Remarkably, the BHs exhibit a horizonless nature for $c_{2}$ values higher than $1.46,1.65$ and $1.03$ corresponding to $\omega=0,1/3,-2/3$ respectively with respect to $\beta=-11.8$ as depicted in lower row of Fig 1.
Refer to caption     Refer to caption    Refer to caption
Figure 2: Behavior of y=2M/r1y=2M/r-1 (solid line) and $y=c_{2}\,r^{-n}$ (red dashed line) as functions of $r$. These plots are presented to examine the maximum number of positive real roots of equation $A(r)=0$ of $f(R,T)$ BHs for $\omega=0,\,1/3,\,-2/3$ cases considering $M=1$. The black and red dots are two intersecting points of the solid curve with the red dashed line providing the locations of the inner horizon $r_{ih}$ and the event horizon $r_{eh}$ respectively.
Furthermore, we adopt a standard graphical method (Fig. 2) to analyze the number of positive real roots of the equation $A(r)=0$ [124, 125], in which the behavior of both sides of the equation $2M/r-1=c_{2}\,r^{-n}$ obtained from equation $A(r)=0$, where $n=[8(\beta\omega+\pi(6\omega+2))]/[16\pi-\beta(\omega-3)]$, are examined as functions of the radial coordinate $r$. The left-hand side (LHS), 2M/r12M/r-1, is a monotonically decreasing function for $r>0$, diverging to positive infinity as $r\to 0^{+}$ and asymptotically approaching 1-1 as $r\to\infty$. In contrast, although the right-hand side (RHS), $c_{2}\,r^{-n}$ is also a monotonically decreasing function, continues asymptotically to zero as $r\to\infty$ for $c_{2}>0$, $n>0$, and diverges to infinity without bound as $r\to 0^{+}$. The points of intersection of these two graphs correspond to the real positive solutions of the equation $A(r)=0$. Depending on the values of $c_{2}$ and $n$, the system can exhibit the following three distinct scenarios:
  • (i)
    Two positive real roots, when the LHS and RHS curves intersect at two distinct points, indicating two physically relevant solutions.
  • (ii)
    A single positive root, when the curves meet at a single point, marking a critical case.
  • (iii)
    No real roots, when the LHS lies entirely below the RHS for all $r>0$, indicating the absence of any intersection.
Fig. 2 demonstrates a case that the equation $A(r)=0$ can admit a maximum of two real positive solutions, though the actual number may reduce to one or zero depending on the parameter values as shown in Fig. 1. Hence, our BHs can have a maximum two number of horizons only. It can be inferred that the spacetime geometry of the considered BHs differs fundamentally from that of the Schwarzschild BH due to the presence of the Cauchy horizon. Thus observing such a horizon structure scenario, we can remark that the structure of a minimally coupled $f(R,T)$ gravity BHs changes typically and reflects a significant change in the spacetime’s causal structure as the model parameters vary.